# Ctrl-GenAug: Controllable Generative Augmentation for Medical Sequence Classification

### v1

@@ IV-B Generative Augmentation-based Sequence Classification
Table II compares the average performance (i.e., accuracy and AUROC) on different training paradigms across a wide range of sequence classifiers, including five CNN-based classic models, three transformer-based models, and three CNN-Transformer hybrid-based methods. It is evident that the performance ceiling of existing popular models, from simple CNN-based to complex CNN-Transformer hybrids, can be further elevated using synthetic samples obtained from our proposed framework, under Real-finetune and Joint-train paradigms. On the carotid dataset, the accuracy/AUROC can be most boosted by 5.44%/0.077 and 6.80%/0.086 through Real-finetune and Joint-train paradigms, respectively. Due to high intra-class variations in TUSC, the baseline exhibits visibly inferior results, with an average accuracy of 62.42% compared to other datasets. However, it can be seen from Table II that samples synthesized by Ctrl-GenAug make significant contributions and enhance the classification performance to an acceptance level. Among all the classifiers, CSN harvests the largest performance gain, which is non-trivially enhanced by 5.41%/0.061 and 3.61%/0.040 accuracy/AUROC under Real-finetune and Joint-train paradigms, respectively. Notably, for ACDC, the overall performance improvement remains significant even with a high baseline. For instance, CSG-3DCT is boosted from 87.16$\rightarrow$88.63(+1.47%) accuracy and from 0.797$\rightarrow$0.820 (+0.023) AUROC under the Joint-train practice. Extensive experimental results show that synthetic sample sets produced by Ctrl-GenAug can be powerful data engines to enhance various diagnostic tasks in medical sequence analysis.
IV-C Improved Diagnosis in Underrepresented High-risk Sets
@@
IV-D Classification Robustness in Out-domain Conditions
We further explored the impact of synthetic samples from Ctrl-GenAug on diagnostic robustness in out-domain settings using the carotid dataset. Training data of the sequence generator consisted of labeled ID and unlabeled OD subsets. We added the hospital identifier in the descriptive text of the ID data to condition the domain distribution. For the OD data, which did not contain the diagnostic class label, we padded the corresponding conditioning vector with zeros, while solely preserving hospital IDs to form descriptive texts. As shown in Table III, the average accuracy and AUROC of three classifiers display better results on the Real-finetune and Joint-train paradigms compared to the Baseline. It validates that diagnostic robustness can be enhanced with the aid of Ctrl-GenAug in scenarios where we only have access to unlabeled cases from additional medical centers due to limited resources.
@@ IV-D Classification Robustness in Out-domain Conditions
IV-E Analysis of Multimodal Conditions Guidance
To illustrate the role of different conditional guidance, we conduct comparative experiments with generators that resort to various banks of conditions for training and sampling. Table IV validates that all proposed conditions are effective for enhancing synthesis and diagnosis tasks. Specifically, without text control, the downstream diagnostic performance is comparatively poor. Fig. 5(b-c) compares typical synthetic thyroid nodule sequences generated under condition banks with and without text guidance. The former is visibly high-fidelity and faithful to the given text prompt (e.g., smooth margin), while the latter exhibits less distinguishable features for diagnosis (e.g., blurry margin). That is to say, the proposed Ctrl-GenAug can create more diagnosis-reliable samples by enhancing semantic steerability in the generation process. Besides, by incorporating image prior knowledge in the sampling process, the average accuracy and AUROC of three classifiers are improved by 4.20% and 0.040, respectively. This proves the domain gap between synthetic and real samples is mitigated by introducing the image prior. Moreover, as shown in Table IV, conditioning the generator on the motion field produces more content-consistent and smoother samples, leading to continuous improvements in diagnostic performance. It can also be observed that the FVD results show a limited correlation with downstream evaluation metrics, confirming the finding in [7].
TABLE IV: Ablation study for different conditional controls in TUSC [69], including class label (C), text (T), image prior (I), and motion field (MF). SlowFast [81] (C1), CSN [82] (C2), and FTC [78] (C3) trained on Joint-train paradigm were used for downstream diagnosis. “Hybrid”, the classifier with CNN-Transformer design.
@@ IV-E Analysis of Multimodal Conditions Guidance
TUSC [69] 4.18 90.40% 95.03% C1 CNN $\downarrow$ 1.80% $\downarrow$ 0.015
@@
IV-F Impact of the Sequential Augmentation Module
To verify the impact of the proposed sequential augmentation module, as shown in Table V, we compared the performance of three generator variants on the synthesis and downstream tasks. Ours-S, Ours-SK, and Ours-SKM denote our ablation studies, including gradually adding sequential attention (’-S’), key-frame/slice attention (’-K’), and motion field attention (’-M’) to the generator with spatial inflation only. Table V shows that the downstream accuracy and AUROC improvements of Ours-SKM are significantly higher than those of Ours-S and Ours-SK in both datasets. In terms of quantitative assessment of sequential coherence, Ours-SKM achieves better results overall than other variants. These prove that this module can help the generator synthesize cross-frame/slice consistent and dynamic-smooth clips, which are diagnosis-promotive for downstream diagnostic tasks. Moreover, it can be observed that the FVD results across these three generators show no significant difference, revealing its limited value for evaluating diagnosis-oriented medical sequence synthesis tasks.
TABLE V: Ablation study for the sequential augmentation module of the generator in TUSC [69] and ACDC [70]. SlowFast [81] (C1), CSN [82] (C2), and FTC [78] (C3) trained on Joint-train paradigm were used for downstream diagnosis.
@@ IV-F Impact of the Sequential Augmentation Module
TUSC [69] Ours-S 3.76 87.44% 91.07% C1 CNN $\downarrow$ 0.36% $\downarrow$ 0.009
@@
ACDC [70] Ours-S 8.75 70.11% 87.62% C1 CNN $\uparrow$ 0.73% $\uparrow$ 0.015
@@
Generative Training Paradigm Augmentation Method Carotid TUSC [69] ACDC [70]
@@
IV-G Effectiveness of the Noisy Synthetic Data Filter
IV-H Impacts of Traditional and Generative Augmentations
@@
We tested the contributions of different data augmentations in three downstream diagnostic tasks. The traditional augmentations used in this experiment can be found in Table I. We report the performance gains of the two variants with data augmentations compared to the augmentation-free setup. For experimental fairness, we kept the training datasets for the None and TraAug variants the same size as the method combining traditional and generative augmentations. As presented in Table VI, the diagnostic performance shows noticeable improvements using the traditional augmentation method. Further coupled with the generative augmentation, the results show more substantial improvements than TraAug. We also observe that in the case of using traditional augmentation, adding extra real training samples yields limited performance gains, even a drop for ACDC. These prove that generative augmentation can effectively supplement traditional data augmentation and provide intrinsically diverse data characteristics for enhancing deep classifier learning, which support the viewpoints of [7].
V Conclusion and discussion
@@
In this study, we present a new and general diffusion-based generative augmentation framework, named Ctrl-GenAug, to facilitate medical sequence classification by leveraging customized and diagnosis-reliable synthetic sequences. To the best of our knowledge, this is the first comprehensive study to investigate the impact of the controllable generative augmentation scheme on aiding medical sequence classification. To improve the quality of synthetic data for promoting downstream classification, we propose a multimodal conditions-guided medical sequence generator that ensures flexibly controllable synthesis across semantic, sequential, and data distribution aspects. Moreover, we propose a highly effective noisy synthetic data filter to better learn from synthetic data, including adaptively filtering diagnosis-inhibitive synthetic sequences at class semantics and sequential levels. This can better connect the synthesis task and the downstream one, thus further enhancing our sequence classification performance. Extensive experiments on 3 medical datasets across multiple organs and modalities, with 11 popular networks trained on 3 paradigms, show the effectiveness and generality of Ctrl-GenAug. Furthermore, we conduct extensive empirical analysis demonstrating that Ctrl-GenAug can be effectively leveraged to improve diagnostic performance in underrepresented high-risk populations and out-domain robustness. We believe that Ctrl-GenAug can serve as a robust and practical data augmentation tool for various clinical scenarios.
@@
We then discuss the limitations of Ctrl-GenAug. First, the denoising diffusion implicit model (DDIM) sampling process is time-consuming due to the large inference steps. This limits the clinical practicality of Ctrl-GenAug for fast and large-scale production. Second, in out-domain conditions, the proposed framework requires target domain data during generator training, which can be tough to obtain in advance in clinical practice. In future work, we will adopt fast-sampling strategies (e.g., AMED-Solver [88]) to better balance the trade-off between sampling time and sample quality. Besides, we will incorporate test-time adaptation methods [89, 90] to provide a more user-friendly tool for out-domain applications in clinical settings. Last, we will attempt to further extend the proposed method to segmentation and detection tasks.
@@
Figure S1: Additional synthetic results on Carotid (a-d), TUSC [69] (e-h), and ACDC [70] (i-l). We used the multimodal conditions joint training strategy [76] during training and flexibly dropped several during sampling.
@@
Shang et al. [49] disease class,
@@ 2 Fundus image
Farooq et al. [50] text classification CNN-based,
@@ 2 Dermoscopy image
Zhang et al. [51] text, edge segmentation CNN-Transformer hybrid-based 1 US, CT, MRI image
Sagers et al. [91] text classification CNN-based 1 Dermoscopy image
Khosravi et al. [92] attribute, 14 pathology labels classification CNN-based 1 X-ray image
Yu et al. [93] pathology labels,
@@
Ktena et al. [25] disease class, attribute classification CNN-based 1 Histopathology,
@@
Luo et al. [7] disease class classification CNN-based 10 Fundus,
@@
Li et al. [24] multi-scale latent
@@
Gong et al. [94] mask segmentation CNN-based 1 MRI volume
Peng et al. [95] mask segmentation CNN-based, CNN-Transformer hybrid-based 3 CT volume
Kebaili et al. [96] tumor attribute, mask segmentation CNN-based 1 MRI volume
Yu et al. [97] mask segmentation CNN-based 1 CT volume
Chen et al. [98] mask segmentation CNN-based,
@@ 3 CT volume
Dorjsembe et al. [99] mask segmentation CNN-based 1 MRI volume
Huang et al. [100] mask segmentation CNN-based 2 OCT volume

### v2

@@ 4.2 Generative Augmentation-based Sequence Classification
Table 2 compares the average performance (i.e., accuracy and AUROC) on different training paradigms across a wide range of sequence classifiers, including five CNN-based classic models, three transformer-based models, and three CNN-Transformer hybrid-based methods. It is evident that the performance ceiling of existing popular models, from simple CNN-based to complex CNN-Transformer hybrids, can be further elevated using synthetic samples obtained from our proposed framework, under Real-finetune and Joint-train paradigms. On the carotid dataset, the accuracy/AUROC can be most boosted by 5.44%/0.077 and 6.80%/0.086 through Real-finetune and Joint-train paradigms, respectively. Due to high intra-class variations in TUSC, the baseline exhibits visibly inferior results, with an average accuracy of 62.42% compared to other datasets. However, it can be seen from Table 2 that samples synthesized by Ctrl-GenAug make significant contributions and enhance the classification performance to an acceptance level. Among all the classifiers, CSN harvests the largest performance gain, which is non-trivially enhanced by 5.41%/0.061 and 3.61%/0.040 accuracy/AUROC under Real-finetune and Joint-train paradigms, respectively. Notably, for ACDC, the overall performance improvement remains significant even with a high baseline. For instance, CSG-3DCT is boosted from 87.16$\rightarrow$88.63(+1.47%) accuracy and from 0.797$\rightarrow$0.820 (+0.023) AUROC under the Joint-train practice. Extensive experimental results show that synthetic sample sets produced by Ctrl-GenAug can be powerful data engines to enhance various diagnostic tasks in medical sequence analysis.
4.3 Improved Diagnosis in Underrepresented High-risk Sets
@@
4.4 Classification Robustness in Out-domain Conditions
We further explored the impact of synthetic samples from Ctrl-GenAug on diagnostic robustness in out-domain settings using the carotid dataset. Training data of the sequence generator consisted of labeled ID and unlabeled OD subsets. We added the hospital identifier in the descriptive text of the ID data to condition the domain distribution. For the OD data, which did not contain the diagnostic class label, we padded the corresponding conditioning vector with zeros, while solely preserving hospital IDs to form descriptive texts. As shown in Table 3, the average accuracy and AUROC of three classifiers display better results on the Real-finetune and Joint-train paradigms compared to the Baseline. It validates that diagnostic robustness can be enhanced with the aid of Ctrl-GenAug in scenarios where we only have access to unlabeled cases from additional medical centers due to limited resources.
@@ 4.4 Classification Robustness in Out-domain Conditions
Table 4: Ablation study for different conditional controls in TUSC (Stanford-AIMI, 2021), including class label (C), text (T), image prior (I), and motion field (MF). SlowFast (Feichtenhofer et al., 2019) (C1), CSN (Tran et al., 2019) (C2), and FTC (Ahmadi et al., 2023) (C3) trained on Joint-train paradigm were used for downstream diagnosis. “Hybrid”, the classifier with CNN-Transformer design.
@@
TUSC (Stanford-AIMI, 2021) 4.18 90.40% 95.03% C1 CNN $\downarrow$ 1.80% $\downarrow$ 0.015
@@
4.5 Analysis of Multimodal Conditions Guidance
To illustrate the role of different conditional guidance, we conduct comparative experiments with generators that resort to various banks of conditions for training and sampling. Table 4 validates that all proposed conditions are effective for enhancing synthesis and diagnosis tasks. Specifically, without text control, the downstream diagnostic performance is comparatively poor. Fig. 6(b-c) compares typical synthetic thyroid nodule sequences generated under condition banks with and without text guidance. The former is visibly high-fidelity and faithful to the given text prompt (e.g., smooth margin), while the latter exhibits less distinguishable features for diagnosis (e.g., blurry margin). That is to say, the proposed Ctrl-GenAug can create more diagnosis-reliable samples by enhancing semantic steerability in the generation process. Besides, by incorporating image prior knowledge in the sampling process, the average accuracy and AUROC of three classifiers are improved by 4.20% and 0.040, respectively. This proves the domain gap between synthetic and real samples is mitigated by introducing the image prior. Moreover, as shown in Table 4, conditioning the generator on the motion field produces more content-consistent and smoother samples, leading to continuous improvements in diagnostic performance. It can also be observed that the FVD results show a limited correlation with downstream evaluation metrics, confirming the finding in (Luo et al., 2024).
Table 5: Ablation study for the sequential augmentation module of the generator in TUSC (Stanford-AIMI, 2021) and ACDC (Bernard et al., 2018). SlowFast (Feichtenhofer et al., 2019) (C1), CSN (Tran et al., 2019) (C2), and FTC (Ahmadi et al., 2023) (C3) trained on Joint-train paradigm were used for downstream diagnosis.
@@ 4.5 Analysis of Multimodal Conditions Guidance
TUSC (Stanford-AIMI, 2021) Ours-S 3.76 87.44% 91.07% C1 CNN $\downarrow$ 0.36% $\downarrow$ 0.009
@@
ACDC (Bernard et al., 2018) Ours-S 8.75 70.11% 87.62% C1 CNN $\uparrow$ 0.73% $\uparrow$ 0.015
@@
4.6 Impact of the Sequential Augmentation Module
To verify the impact of the proposed sequential augmentation module, as shown in Table 5, we compared the performance of three generator variants on the synthesis and downstream tasks. Ours-S, Ours-SK, and Ours-SKM denote our ablation studies, including gradually adding sequential attention (’-S’), key-frame/slice attention (’-K’), and motion field attention (’-M’) to the generator with spatial inflation only. Table 5 shows that the downstream accuracy and AUROC improvements of Ours-SKM are significantly higher than those of Ours-S and Ours-SK in both datasets. In terms of quantitative assessment of sequential coherence, Ours-SKM achieves better results overall than other variants. These prove that this module can help the generator synthesize cross-frame/slice consistent and dynamic-smooth clips, which are diagnosis-promotive for downstream diagnostic tasks. Moreover, it can be observed that the FVD results across these three generators show no significant difference, revealing its limited value for evaluating diagnosis-oriented medical sequence synthesis tasks.
4.7 Effectiveness of the Noisy Synthetic Data Filter
@@
Generative Training Paradigm Augmentation Method Carotid TUSC (Stanford-AIMI, 2021) ACDC (Bernard et al., 2018)
@@
4.8 Impacts of Traditional and Generative Augmentations
@@
We tested the contributions of different data augmentations in three downstream diagnostic tasks. The traditional augmentations used in this experiment can be found in Table 1. We report the performance gains of the two variants with data augmentations compared to the augmentation-free setup. For experimental fairness, we kept the training datasets for the None and TraAug variants the same size as the method combining traditional and generative augmentations. As presented in Table 6, the diagnostic performance shows noticeable improvements using the traditional augmentation method. Further coupled with the generative augmentation, the results show more substantial improvements than TraAug. We also observe that in the case of using traditional augmentation, adding extra real training samples yields limited performance gains, even a drop for ACDC. These prove that generative augmentation can effectively supplement traditional data augmentation and provide intrinsically diverse data characteristics for enhancing deep classifier learning, which support the viewpoints of (Luo et al., 2024).
5 Conclusion and discussion
In this study, we present a new and general diffusion-based generative augmentation framework, named Ctrl-GenAug, to facilitate medical sequence classification by leveraging customized and diagnosis-reliable synthetic sequences. To the best of our knowledge, this is the first comprehensive study to investigate the impact of the controllable generative augmentation scheme on aiding medical sequence classification.
@@ 5 Conclusion and discussion
To improve the quality of synthetic data for promoting downstream classification, we propose a multimodal conditions-guided medical sequence generator that ensures flexibly controllable synthesis across semantic, sequential, and data distribution aspects. Moreover, we propose a highly effective noisy synthetic data filter to better learn from synthetic data, including adaptively filtering diagnosis-inhibitive synthetic sequences at class semantics and sequential levels. This can better connect the synthesis task and the downstream one, thus further enhancing our sequence classification performance. Extensive experiments on 3 datasets across multiple organs and modalities, with 11 popular networks trained on 3 paradigms, show the effectiveness and generality of Ctrl-GenAug. Furthermore, the empirical analysis validates that Ctrl-GenAug can be effectively leveraged to improve diagnostic performance in underrepresented high-risk populations and out-domain robustness. We believe that Ctrl-GenAug can serve as a robust and practical data augmentation tool for various clinical scenarios.
@@
We then discuss the limitations of Ctrl-GenAug. First, the denoising diffusion implicit model (DDIM) sampling process is time-consuming due to the large inference steps. This limits the clinical practicality of Ctrl-GenAug for fast and large-scale production. Second, in out-domain conditions, the proposed framework requires target domain data during generator training, which can be tough to obtain in advance in clinical practice. In future work, we will adopt fast-sampling strategies (e.g., AMED-Solver (Zhou et al., 2024b)) to better balance the trade-off between sampling time and sample quality. Besides, we will incorporate test-time adaptation methods (Huang et al., 2023b; Luo et al., 2023; Huang et al., 2023c; Zhang et al., 2024b) to provide a more user-friendly tool for out-domain applications in clinical settings. Last, we will attempt to further extend the proposed method to segmentation and detection tasks (Huang et al., 2025; Tao et al., 2025).
@@
Figure S1: Additional synthetic results on Carotid (a-d), TUSC (Stanford-AIMI, 2021) (e-h), and ACDC (Bernard et al., 2018) (i-l). We used the multimodal conditions joint training strategy (Huang et al., 2023a) during training and flexibly dropped several during sampling.
@@
(Shang et al., 2023) disease class,
@@ 2 Fundus image
(Farooq et al., 2024) text classification CNN-based,
@@ 2 Dermoscopy image
(Zhang et al., 2024c) text, edge segmentation CNN-Transformer hybrid-based 1 US, CT, MRI image
(Sagers et al., 2023) text classification CNN-based 1 Dermoscopy image
(Khosravi et al., 2024) attribute, 14 pathology labels classification CNN-based 1 X-ray image
(Yu et al., 2024a) pathology labels,
@@
(Ktena et al., 2024) disease class, attribute classification CNN-based 1 Histopathology,
@@
(Luo et al., 2024) disease class classification CNN-based 10 Fundus,
@@
(Li et al., 2024a) multi-scale latent
@@
(Gong et al., 2024) mask segmentation CNN-based 1 MRI volume
(Peng et al., 2024) mask segmentation CNN-based, CNN-Transformer hybrid-based 3 CT volume
(Kebaili et al., 2024) tumor attribute, mask segmentation CNN-based 1 MRI volume
(Yu et al., 2024b) mask segmentation CNN-based 1 CT volume
(Chen et al., 2024) mask segmentation CNN-based,
@@ 3 CT volume
(Dorjsembe et al., 2024a) mask segmentation CNN-based 1 MRI volume
(Huang et al., 2024a) mask segmentation CNN-based 2 OCT volume